Grade 12Physical Chemistry

A compound has formula AB3 type ,A occupies lattice in cubic close packing what is the coordination number of A and B

Nizam
6 Years agoGrade 12

1 Answer

Vikas TU
6 Years ago
Dear stuent 
A atoms are at eight corners of the cube.
Therefore,the number of A atoms in the unit cell =8/8=1
B atoms are at the face-centre of six faces.
Therefore,its share in the unit cell =6/2=3
∴ The formula is AB3